Triangle ABC is isosceles with AB congruent to AC. Extend segment BA to a point T (in other wordds, A should be between B and T.) Prove that angle TAC must be twice the size of angle ABC. Angle TAC is called one of the exterior angles of triangle ABC.
Answer by Nate(3500)

Angles CBA and BCA will be known as "x", then:
BAC = 180 - 2x
TAC = 180 - BAC
TAC = 180 - ( 180 - 2x )
TAC = 2x while ABC = x